OpenStudy (anonymous):

How do you find the product of: 4g(g-3)-2g(g-2)=2(g2+3)-6g-2?

OpenStudy (anonymous):

Let 4g(g-3)-2g(g-2)=2(g^2+3)-6g-2, becomes:4g^2-12g-2g^2-4g=2g^2+6-6g-2 4g^2-2g^2-2g^2-12g-4g+6g=6-2 -10g=4 g=-(4/10) g=-2/5.

OpenStudy (anonymous):

I disagree Calc. It comes out to be -2.

OpenStudy (anonymous):

show your work